OpenStudy (anonymous):

A 1,600 kg train car rolling freely on level track at 16 m/s bumps into a 1.0 × 103 kg train car moving at 10.0 m/s in the same direction, and the two latch onto each other and continue together. What is their final speed? 16 m/s 6.0 m/s 10 m/s 14 m/s

OpenStudy (jamesj):

Again, use Conversation of Momentum. The momentum of the two trains before bumping into each other is the same as it is afterwards. Can you write down the momentum before they bump?

OpenStudy (anonymous):

whats the formula again?

OpenStudy (anonymous):

is the ans 16m/s

OpenStudy (anonymous):

thnx....jamesj left me hangin

OpenStudy (anonymous):

jamesj wants you to try to understand it yourself

OpenStudy (anonymous):

i know but I asked for the formula n he left

OpenStudy (anonymous):

If i knew the formula Id do it on my own...he was helping me...n I appreciate it....jus need more help still

OpenStudy (anonymous):

total momentum before collision=total momentum after collision

OpenStudy (anonymous):

thnx...again

OpenStudy (anonymous):

Correction=the ans is 14m/s not 16 Didn't notice the power '1x103' write it like 1x10^3

OpenStudy (anonymous):

yea...but I didnt wanna say you were wrong...still helped though :)

OpenStudy (jamesj):

@physicsme, don't give the answer. Give the formula. Momentum = (mass) x (velocity)

OpenStudy (anonymous):

thats all i really wated was the formula...

OpenStudy (anonymous):

wanted*

OpenStudy (anonymous):

okay, actually I was practicing different questions so just wanted to solve it and check the answer.

OpenStudy (jamesj):

Hence, for example the momentum of the first train before the collision is p1 = m1.v1 = (1600 kg)(16 m/s) = 25,600 kg.m/s The momentum of the second train is p2 = m2.v2 = (103 kg)(10 m/s) = 1,030 kg.m/s [btw, is that mass of the train car right? Looks wrong. If so, fix this calculation] Now after the collision, they have the same velocity v' and their combined mass is m = m1 + m2 Now set the sum of the momentum before = momentum afterwards